3. A graduate student believes that his starting salary will be based on his thesis professors salary, in the form of the equation Y 400000.60 (X -10000) where Y is the students average salary and X is the thesis professors salary. Consider a population of 100 graduate students, with an average thesis professor salary of $125,000 and standard deviation $30000 (a) Find the expected student salary based on this equation (b) Find and interpret the standard deviation of student salary.(c) Would it be unexpected to see a student salary of $200,000? Explain. Hint: How many standard deviations is that away from the average?
Answer #1

(a)

The given equation is written as follows:

y = 40000 0.60X-(0.60 × 10000)

i.e.,

0.60X + 34000 (1)

By Theorem:

E(aX + b) = aE(X) + b (2)

Applying (2), equation (1) becomes:

E(Y)-0.60E(X) + 34000 (3)

Given:
E(X)=125000 (4)

Substituting (4), equation (3) becomes:

E(Y)-(0.60 × 125000) + 34000-109000

So,

Answer is:

109000

(b)

By Theorem:

Var(aX + b) = a-Var(X) (5)

Given:

Var(X) 300002 (6)

Substituting (6), equation (5) becomes:

Var(Y) 0.602Var(X)-0.36 × 300002-324.000.000

So,

Answer is:

324,000,000

(c)

Standard deviation of Y is given by:

V324,000, 000 18, 000

(200,000 - 109,000)/18,000 = 5.0556

It would be unusual to see a student salary of $200,000 because 200,000 is 5.0556 > 2 standard deviations away from the average.